### v1

@@ IV-B Performance Comparison by Architecture
Table II summarizes Elo ratings of representative models. Classical CNN engines achieved the highest rating of 1891, followed by hybrid CQCNN models with QNNs, then those with QCNNs, and finally standalone QCNN engines.
Models using 18 qubits tended to perform worse than those with fewer qubits, likely due to overparameterization relative to the simplicity of tic-tac-toe. The two CNN engines listed give a rough idea of the position of the CQCNN engines.
TABLE II: Main engines and the Elo rating: Roughly speaking, the engines are divided into the following order from highest rating to lowest: classical, CQCNN with QNN as the hidden layer, CQCNN with QCNN as the hidden layer, and QCNN. Even QCNN, which has the lowest rating among the engines, has a difference of 1609, or nearly 100, which gives it a win rate of about 65%, referring to Fig. 5.
@@
classical neural networks stronger 1891
Sampler 7 qubit HEE & EfficientSU2 1809
Estimator 7 qubit HEE & RealAmplitudes 1776
Estimator 18 qubit TPE & RealAmplitudes 1762
Sampler 18 qubit TPE & EfficientSU2 1752
cqcqcnn 10 qubit ZFeatureMap 1751
cqcqcnn 18 qubit HEE 1713
classical neural network weaker 1712
qcnn 18 qubit ZZFeatureMap 1609
random policy - 1500
@@ IV-C Demonstration Using IBM Quantum Hardware
To validate the feasibility of executing the proposed engine on real quantum hardware, we deployed the trained model on the ibm_kawasaki backend via the IBM Quantum platform. Two matches were conducted: one where the quantum engine played first and another where it played second.
@@
Figure 10: Quantum computer vs Human: The quantum computer was first and a human was second. Looking at the quantum computer’s moves, the first move was not so bad, but in the middle of the game it could be seen that it was trying to align vertically. However, the game was over without defence when the human tried to align diagonally in the endgame.
Figure 11: Human vs Quantum computer: A human was first and the quantum computer was second. The quantum computer’s move showed that it was trying to align vertically, but it did not interfere with the human move and the game was over.
@@ V The Overhead of Engines through QI Model
V-A Toy Model Settings
Figure 12: Quantum communication model: In Model 1, the client embeds classical data into a qubit, passes it to the server, which calculates the ansatz and then returns it to the client. The client measures the qubit and obtains an output. In Model 2, the client embeds classical data into qubits, passes them to the server, calculates the ansatz, and then makes a measurement. The measurement results are returned to the client and an output is obtained.
We consider two models of quantum communication channels, depending on the quantum resources available on the client side.
@@ V-A Toy Model Settings
  • Model 2: The client cannot measure the quantum states after server-side computation; instead, the server performs the measurement and returns the result to the client.
@@
We employ a 7-qubit CQCNN using Sampler, with different Embedding and Ansatz. Noise is introduced by inserting a rotation gate between the quantum circuits of the engine, with its parameters randomized to pseudo-realize the noise effect.
@@
$\sigma(d)=10^{\frac{\alpha}{10}d}-1$ (4)
@@
    Fixed Distance Experiment: We assess the engine performance under three noise patterns, with the distance set to $15\,\mathrm{km}$.
@@
    • Pattern B: Noiseless during training, but noised during evaluation.
@@
    Distance-Dependent Experiment: We vary the communication distance from $0\,\mathrm{km}$ to $15\,\mathrm{km}$ in increments of $3\,\mathrm{km}$ and evaluate the engine under Pattern B.
@@ V-B The Effect of Noisy Quantum Communication
Table. III presents the results for different noise models and engines. Fig. 13 and Fig. 14 illustrate the rating transitions of the engines listed in the first row of Table. III.
TABLE III: The result of using different noise models and engines: The table features 3 different engines in each model.
engine model train evaluate rating
Sampler HEE EfficientSU2 1 noiseless noiseless 1728
noiseless noised 1630
noised noised 1619
2 noiseless noiseless 1692
noiseless noised 1621
noised noised 1537
Sampler ZFeatureMap RealAmplitudes 1 noiseless noiseless 1685
noiseless noised 1684
noised noised 1573
2 noiseless noiseless 1711
noiseless noised 1663
noised noised 1554
Sampler ZZFeatureMap RealAmplitudes 1 noiseless noiseless 1731
noiseless noised 1675
noised noised 1431
2 noiseless noiseless 1677
noiseless noised 1651
noised noised 1548
The Elo ratings of all engines in Table. III follow the order Pattern A, Pattern B and Pattern C, indicating that noise negatively impacts performance. However, the extent of the rating drop varies depending on the engine.
For instance, in Model 1, the engine using HEE and EfficientSU2 shows a significant gap between Pattern A and the others, whereas Patterns B and C yield similar ratings. In contrast, the engine using ZFeatureMap and RealAmplitudes in Model 1 exhibits similar ratings for Patterns A and B, but a significant drop in Pattern C. These differences suggest that the robustness to noise depends on the choice of embedding and ansatz.
Figure 13: Transition of ratings when evaluating engines using HEE and EfficientSU2: he_1 represents Model 1 using HEE and EfficientSU2. he_1_f_f is the Pattern A, he_1_f_t is the Pattern B, and he_1_t_t is the Pattern C.
Figure 14: Transition of ratings when evaluating engines using HEE and EfficientSU2: he_2 represents Model 2 using HEE and EfficientSU2. he_2_f_f is the Pattern A, he_2_f_t is the Pattern B, and he_2_t_t is the Pattern C.
Fig. 13 and Fig. 14 show the progression of the rating in every 100 games. In both figures, Pattern A exhibits a steady increase in ratings, while Patterns B and C show slower growth due to the effects of noise.
Figure 15: Transition of ratings per distance: he_model1 represents Model 1 using HEE and EfficientSU2, and he_model2 is Model 2 using HEE and EfficientSU2.
Fig. 15 illustrates the rating transition as a function of distance. As shown in the figure, there is a sharp drop in the ratings between 00 and $3\,\mathrm{km}$, followed by a more gradual decline as the distance increases. This suggests that the initial increase in noise significantly impacts accuracy, but beyond a certain distance, the degradation stabilizes.
@@ VI Conclusion
In this study, we built game engines using classical, quantum, and classical-quantum hybrid neural networks and compared their performances. By comparing the performance of each engine in terms of ratings, we can determine the relative strength of the engines. In the same way that the development of computers and algorithms has led to the development of classical game engines such as chess, the development of quantum computers and quantum algorithms is expected to contribute to the development of quantum game engines as they are developed.
@@
Although it was difficult for the quantum computer to outperform classical methods due to the game of tic-tac-toe in this study, the objective of comparing the different engines through the rating was achieved. We also modeled a simple noisy communication channel, which can be used not only for the game solver task but also as a guide for developing various quantum network applications, and verified the amount of overhead observed. This is a necessary indicator not only for simple applications but also for the development of hidden applications. It will be a stepping stone to develop use cases of quantum network applications that are better than classical ones, while taking noise into account.
Appendix A Quantum Engine Configurations
A-A Quantum Circuits
Fig. 16 shows Embedding and Ansatz used in this paper.
@@ A-A Quantum Circuits
ZFeatureMap, ZZFeatureMap, RealAmplitudes and EfficientSU2 are reffered to the qiskit library. HEE and TPE are reffered to the circuit for embedding in  [23], and QCNN is reffered to the circuit in  [24] and Qiskit machine learning tutorial.
Figure 16: Embedding and Ansatz used in engines: (a) Embedding is a quantum circuit for embedding certain classical data into qubits. ZFeatureMap and ZZFeatureMap were built using the qiskit library, while HEE and TPE were constructed by referring to the circuit for embedding in  [23]. (b) Ansatz is a circuit for processing data embedded in qubits using weight parameters. RealAmplitudes and EfficientSU2 were built using the qiskit library. QCNN referred to the circuit in [24] and Qiskit machine learning tutorial.
A-B Engine Properties and Rating
Table IV lists the 46 engines mentioned in Sec. IV and their ratings and other detailed information.
TABLE IV: Properties and Elo ratings of 46 engines: This table is divided into eight rows, which can be classified as follows: (a) the first row, (b) rows 2-5, (c) rows 6-7, and (d) the eighth row. (a) is 2 types of engines using only CNN. (b) is 32 types of engines using CQCNN with QNN. (c) is 8 types of engines using CQCNN with QCNN. (d) is 4 types of engines using only QCNN.
@@ A-B Engine Properties and Rating
classical neural network stronger 1891 10057 - 10057 - -
weaker 1712 297 - 297 - -
sampler 7 qubit ZFeatureMap+RealAmplitudes 1705 1231 21 1252 6 10
ZZFeatureMap+RealAmplitudes 1625 48 43
HEE+RealAmplitudes 1735 12 15
TPE+RealAmplitudes 1686 6 9
ZFeatureMap+EfficientSU2 1716 35 1266 6 12
ZZFeatureMap+EfficientSU2 1568 48 45
HEE+EfficientSU2 1809 12 17
TPE+EfficientSU2 1779 6 11
sampler 18 qubit ZFeatureMap+RealAmplitudes 1737 2359485 54 2359539 17 21
ZZFeatureMap+RealAmplitudes 1550 323 120
HEE+RealAmplitudes 1535 34 37
TPE+RealAmplitudes 1693 17 20
ZFeatureMap+EfficientSU2 1662 90 2359575 17 23
ZZFeatureMap+EfficientSU2 1476 323 122
HEE+EfficientSU2 1708 34 39
TPE+EfficientSU2 1752 17 22
estimator 7 qubit ZFeatureMap+RealAmplitudes 1725 142 21 163 6 10
ZZFeatureMap+RealAmplitudes 1491 48 43
HEE+RealAmplitudes 1776 12 15
TPE+RealAmplitudes 1686 6 9
ZFeatureMap+EfficientSU2 1720 35 177 6 12
ZZFeatureMap+EfficientSU2 1452 48 45
HEE+EfficientSU2 1683 12 17
TPE+EfficientSU2 1625 6 11
estimator 18 qubit ZFeatureMap+RealAmplitudes 1700 351 54 405 17 21
ZZFeatureMap+RealAmplitudes 1553 323 120
HEE+RealAmplitudes 1742 34 37
TPE+RealAmplitudes 1762 17 20
ZFeatureMap+EfficientSU2 1633 90 441 17 23
ZZFeatureMap+EfficientSU2 1564 323 122
HEE+EfficientSU2 1753 34 39
TPE+EfficientSU2 1729 17 22
cqcqcnn 10 qubit ZFeatureMap 1751 154 55 209 40 17
ZZFeatureMap 1299 130 68
HEE 1602 49 25
TPE 1645 40 16
cqcqcnn 18 qubit ZFeatureMap 1698 270 99 369 72 17
ZZFeatureMap 1578 378 116
HEE 1713 89 33
TPE 1398 72 16
qcnn 18 qubit ZFeatureMap 1585 - 99 99 72 17
ZZFeatureMap 1609 - 378 116
HEE 1493 - 89 33
TPE 1339 - 72 16
Acknowledgment
SK, HK, SN, and TS are supported by JST Moonshot R&D Grant Number JPMJMS226C. SK and TS are also supported by JST COI-NEXT Grant Number JPMJPF2221. SN is also supported by JSPS KAKENHI Grant Number JP2KJ1436. TS is also supported by MEXT KAKENHI Grant Number 22K1978.

### v2

@@ IV-B Performance Comparison by Architecture
Table. II summarizes Elo ratings of representative models. The models listed in the table are the top-performing examples for each architecture; the full results for all 54 engines are detailed in the Appendix B. In this research, the observed fluctuation in Elo rating for a given model was approximately 20-30 points. As shown in Fig. 5, this level of difference corresponds to a variation in win rate of less than 6%. We therefore consider a rating difference of approximately 70 or more to be indicative of a significant performance gap.
TABLE II: Main engines and the Elo rating: The rating is the final result from a round-robin tournament where it was updated every 100 games.
@@
Sampler 16 qubits TPE+RealAmplitudes 1624.44
Sampler 8 qubits TPE+RealAmplitudes 1598.82
Estimator 16 qubits HEE+RealAmplitudes 1578.74
Estimator 8 qubits
ZZFeatureMap
+RealAmplitudes
1554.54
Classical neural network stronger 1546.19
Estimator 18 qubits (only QCNN) HEE+QCNN 1468.43
Estimator 9 qubits (only QNN) TPE+EfficientSU2 1455.61
The two CNN engines listed give a rough idea of the position of the HNN engines. HNN with QNN engines achieved the highest rating of 1624.44, followed by classical CNN engines, then QCNN engines, and QNN engines. We can see that ratings are higher as the number of qubits is increased in these settings. These engines are picked up from different categories, and they differ in the combinations of the circuits, which is the highest.
This rating result can be considered a single example. The ratings for the classical and HNN engines are likely influenced by the construction of the classical layer, the methods of encoding classical information (like the board state) into quantum states, and the process of extracting results via measurement. In this research, we employed general-purpose quantum circuits; therefore, there is likely room for performance improvement by implementing architectures more specialized for tic-tac-toe.
Although the Sampler 16 qubit and Sampler 8 qubit engines achieved the highest ratings, it is believed that their classical layers are a significant contributing factor. In the Sampler method, the classical layer following the quantum layer has a remarkably large number of parameters, which can be thought to increase the engine’s strength. On the other hand, the Estimator 16 qubits and Estimator 8 qubits engines, which have fewer parameters, still recorded higher ratings than the classical neural network. This suggests that the number of parameters is not necessarily the sole determinant of an engine’s strength.
The engines that used only a quantum layer had lower ratings than the classical engines. This indicates that there is room for improvement in areas such as inputting the board state information and processing the output from the neural network. As mentioned earlier, it also highlights the potential for enhancement by implementing quantum circuits more specialized for tic-tac-toe.
Figure 10: Rating progression of the main engines: Since each engine plays 100 games, ratings are calculated every 100 games. Each curve corresponds to the engines in Table. II.
Fig. 10 shows the rating progression of the main engines. We can see the engines appear to be broadly divided into two groups: the top five engines, including the classical neural network, and the two engines constructed only with quantum layers. The higher-ranked engines tend to increase their ratings in a relatively stable manner, while the lower-ranked engines show a tendency to stagnate around their initial rating values. Since the rating calculation is based on the number of wins and losses, excluding draws, it can be inferred that a stagnating rating implies a higher number of draws.
@@ IV-C Demonstration Using IBM Quantum Hardware
To validate the feasibility of executing the proposed engine on real quantum hardware, we deployed the trained model the strongest in Table. II on the ibm_torino backend via the IBM Quantum platform. Two matches were conducted: one where the quantum engine played first and another where it played second.
In Fig. 11, the quantum computer was first. The first move was good, but in the middle of the game, it could be seen that it was trying to align diagonally. However, the game was over because, in addition to trying to form a line in the area human already put, without defense when human tried to align horizontally in the endgame.
In Fig. 12, the quantum computer was second. The quantum computer’s move showed that it was trying to align diagonally, but it did not interfere with the human move and the game was over.
@@
Figure 11: Quantum computer vs Human: The quantum computer was first and a human was second.
Figure 12: Human vs Quantum computer: A human was first and the quantum computer was second.
@@ V The Overhead of Engines through QI Model
V-A Two Toy Model Settings
Figure 13: Quantum communication model: In Model 1, the client embeds classical data into a qubit, passes it to the server, which calculates the ansatz and then returns it to the client. The client measures the qubit and obtains an output. In Model 2, the client embeds classical data into qubits, passes them to the server, calculates the ansatz, and then makes a measurement. The measurement outcomes are returned to the client and an output is obtained.
We consider two models of quantum communication channels, depending on the quantum resources availability on the client side.
@@ V-A Two Toy Model Settings
  • Model 2: The client cannot measure the quantum states after server-side computation; instead, the server performs the measurement and returns the outcomes to the client.
@@
We employ a 8-qubit HNN using Estimator, with different Embedding. Noise is introduced by inserting rotation gates between the quantum circuits of the engine, with its parameters randomized to pseudo-realize the noise effect.
@@
$\sigma(d)=10^{\frac{\alpha}{10}d}-1,$ (4)
@@
    Fixed Distance Experiment: We assess the engine performance under three noise patterns, with the distance set to $100\,\mathrm{km}$.
@@
    • Pattern B: Noiseless during training, but noisy during evaluation.
@@
    Distance-Dependent Experiment: We vary the communication distance from $10^{-2}\,\mathrm{km}$ to $10\,\mathrm{km}$ and evaluate the engine under Pattern B.
In this section, we determine the final rating by playing 10,000 games against a random opponent and updating the rating every 100 games.
@@ V-B The Effect of Noisy Quantum Communication
Table. III presents the results for different noise models and engines.
When comparing Pattern A and Pattern B, almost all engines achieved higher ratings with Pattern A than with Pattern B. This indicates that the ratings are susceptible to noise. However, the extent of the rating drop varies depending on the engine.
When comparing Pattern B and Pattern C, some of the engines show that Pattern B yields higher ratings, and others show that Pattern C’s rating is higher or there is almost no change. This suggests that while noise can adversely affect an engine’s performance, in some cases, it can conversely enhance it. We believe this is likely influenced by factors such as the engine’s level of training, for example, the size of the feature space it explores.
The degree of rating fluctuation due to the patterns varies for each engine. While some engines show little difference depending on the pattern, others exhibit moderate fluctuations. Those that fluctuate significantly are thought to be heavily influenced by noise. We believe this is also heavily influenced by the way the quantum circuits are constructed.
TABLE III: The result of using different noise models and engines: The table features three different engines in each model, and the rating is updated every 100 games and finally determined by playing through 10,000 games against a random engine.
engine model noise pattern rating
Estimator ZFeatureMap RealAmplitudes 1 A 1515.26
B 1494.63
C 1535.15
2 A 1523.31
B 1497.44
C 1494.44
Estimator ZZFeatureMap RealAmplitudes 1 A 1533.80
B 1547.76
C 1487.83
2 A 1540.22
B 1522.41
C 1528.41
Estimator TPE RealAmplitudes 1 A 1573.57
B 1536.29
C 1510.10
2 A 1526.79
B 1524.69
C 1527.20
Fig. 14 illustrates the rating transition as a function of distance. As the figure shows, the rating gradually decreases as the distance $d$ increases. For the engines used in this research, when comparing Model 1 and Model 2, the rate of this decrease appears to be roughly the same. This indicates that the magnitude of the noise affects the rating. Then, around $d=10$, the rating tends to level off, indicating that the engine’s sensitivity to noise saturates at longer distances. When comparing Model 1 and Model 2, the rating for Model 2 begins to stagnate at a smaller distance. This can be attributed to the fact that the magnitude of noise affecting Model 2 is half that of Model 1, and therefore, the impact on its rating is also smaller.
Figure 14: Transition of ratings per distance: The type of engine used in this graph is the following; embedding is HEE, the ansatz is RealAmplitudes, and the noise pattern is Pattern B.
Figure 15: Embedding and Ansatz used in engines: (a) Embedding is a quantum circuit for embedding certain classical data into qubits. ZFeatureMap and ZZFeatureMap were built using the qiskit library, while HEE and TPE were constructed by referring to the circuit for embedding in  [Thanasilp_2023]. (b) Ansatz is a circuit for processing data embedded in qubits using weight parameters. RealAmplitudes and EfficientSU2 were built using the qiskit library. QCNN referred to the circuit in [Vatan_2004] and Qiskit machine learning tutorial.
@@ VI Conclusion
In this research, we built game engines using classical, quantum, and classical-quantum hybrid neural networks and compared their performances. We introduced Elo rating to comprehensively evaluate classical and quantum systems, and by comparing the performance of each engine in terms of ratings, we can determine the relative strength of the engines. In the same way that the development of computers and algorithms has led to the development of classical game engines such as chess, the development of quantum computers and quantum algorithms is expected to contribute to the development of quantum game engines as they are developed.
@@
In this paper, the classical-quantum hybrid method achieved better results than the classical method. Although this result is expected to vary depending on how each engine is constructed and the type of game being played, the objective of comparing the different engines through the rating was achieved. We also modeled a simple noisy communication channel, which can be used not only for the game solver task but also as a guide for developing various quantum network applications, and verified the amount of overhead observed. This is a necessary indicator not only for simple applications but also for the development of hidden applications. It will be a stepping stone to develop use cases of quantum network applications that are better than classical ones, while taking noise into account.
Building on this foundation, there are several promising avenues for future research. It would be valuable to develop more comprehensive frameworks that establish unified standards for comparing quantum and classical systems. This includes exploring more complex tasks where neither paradigm holds a clear, pre-existing advantage, thereby providing a more neutral ground for performance evaluation. Furthermore, as quantum hardware continues to scale, investigating these larger competitive environments and refining the corresponding evaluation methodologies will be crucial in the ongoing search for quantum advantage.
Appendix A Quantum Circuits used in the engines
@@
Fig. 15 shows the Embedding and Ansatz used in this paper. ZFeatureMap, ZZFeatureMap, RealAmplitudes and EfficientSU2 are reffered to the qiskit library. HEE (Hardware Efficient Embedding) and TPE (Tensor Product Embedding) are reffered to the circuit for embedding in  [Thanasilp_2023], and QCNN is reffered to the circuit in  [Vatan_2004] and Qiskit machine learning.
Appendix B Engine Properties and Rating
Table IV lists the ratings and detailed properties of all 46 engines constructed in this research. The engines are broadly divided into three types, classical, classical-quantum hybrid, and quantum. Within those categories, they are further subdivided by qubit count, post-measurement processing, and the choice of Embedding and Ansatz methods.
First, for the classical CNN engines (row 1), we created two versions, ”Stronger” and ”Weaker”. The table shows that the difference in their ratings can be attributed to the large difference in their parameter counts.
Next, for the Hybrid NN engines (rows 2-5), their respective ratings do not change significantly with the number of classical parameters. This suggests that the quantum circuit configuration is the dominant contributor to performance, particularly the choice of Embedding and Ansatz. Furthermore, for the tic-tac-toe task, an increase in circuit depth tends to lower the rating. This is likely task-dependent; for a simple task like tic-tac-toe, a shallower circuit may be sufficient, and a larger depth could lead to training difficulties within our experimental setup.
Finally, the QNN and QCNN engines show low overall ratings, which resulted in a rating below the initial 1500. This may be due to issues with the compatibility of this specific Embedding and Ansatz pairing, which warrants further investigation.
TABLE IV: Properties and Elo ratings of 54 engines: This table is divided into seven rows, which can be classified as follows: (a) the first row, (b) rows 2-5, and (c) rows 6-7. (a) is 2 models of engines using only CNN. (b) is 40 models of engines using HNN with QNN or QCNN. (c) is 12 models of engines using only QNN or QCNN. The rating is the final result from a round-robin tournament where it was updated every 100 games, and the number of parameters indicates the number of trainable weights for the classical and quantum layers, respectively. The number of CX gates and the quantum circuit depth are shown for reference.
@@ Appendix B Engine Properties and Rating
classical neural network stronger 1546.19 10057 - 10057 - -
weaker 1388.85 297 - 297 - -
Sampler 8 qubits ZFeatureMap+RealAmplitudes 1418.76 2393 16 2409 7 11
ZZFeatureMap+RealAmplitudes 1559.11 63 50
HEE+RealAmplitudes 1587.29 14 17
TPE+RealAmplitudes 1598.82 7 10
ZFeatureMap+EfficientSU2 1584.18 32 2425 7 13
ZZFeatureMap+EfficientSU2 1575.87 63 52
HEE+EfficientSU2 1556.22 14 19
TPE+EfficientSU2 1580.00 7 12
Sampler 16 qubits ZFeatureMap+RealAmplitudes 1579.19 589913 32 589945 15 19
ZZFeatureMap+RealAmplitudes 1487.97 255 106
HEE+RealAmplitudes 1607.00 30 33
TPE+RealAmplitudes 1624.44 15 18
ZFeatureMap+EfficientSU2 1562.35 64 589977 15 21
ZZFeatureMap+EfficientSU2 1552.84 255 108
HEE+EfficientSU2 1623.85 30 35
TPE+EfficientSU2 1569.38 15 20
Estimator 8 qubits ZFeatureMap+RealAmplitudes 1537.84 161 16 177 7 11
ZZFeatureMap+RealAmplitudes 1554.54 63 50
HEE+RealAmplitudes 1553.05 14 17
TPE+RealAmplitudes 1416.87 7 10
ZFeatureMap+EfficientSU2 1493.48 32 193 7 13
ZZFeatureMap+EfficientSU2 1543.56 63 52
HEE+EfficientSU2 1494.05 14 19
TPE+EfficientSU2 1489.77 7 12
ZFeatureMap+QCNN 1521.20 125 36 161 32 17
ZZFeatureMap+QCNN 1527.72 88 56
HEE+QCNN 1552.64 39 23
TPE+QCNN 1491.45 32 16
Estimator 16 qubits ZFeatureMap+RealAmplitudes 1487.63 313 32 345 15 19
ZZFeatureMap+RealAmplitudes 1506.73 255 106
HEE+RealAmplitudes 1578.74 30 33
TPE+RealAmplitudes 1503.81 15 18
ZFeatureMap+EfficientSU2 1532.11 64 377 15 21
ZZFeatureMap+EfficientSU2 1547.77 255 108
HEE+EfficientSU2 1476.19 30 35
TPE+EfficientSU2 1495.04 15 20
ZFeatureMap+QCNN 1453.57 241 72 313 64 17
ZZFeatureMap+QCNN 1482.25 304 104
HEE+QCNN 1496.90 79 31
TPE+QCNN 1518.80 64 16
Estimator 9 qubits
(only QNN)
ZFeatureMap+RealAmplitudes 1391.46 - 18 18 8 12
ZZFeatureMap+RealAmplitudes 1374.80 - 80 57
HEE+RealAmplitudes 1319.18 - 16 19
TPE+RealAmplitudes 1313.40 - 8 11
ZFeatureMap+EfficientSU2 1348.71 - 36 36 8 14
ZZFeatureMap+EfficientSU2 1425.78 - 80 59
HEE+EfficientSU2 1393.57 - 16 21
TPE+EfficientSU2 1455.61 - 8 13
Estimator 18 qubits
(only QCNN)
ZFeatureMap+QCNN 1419.14 - 81 81 72 17
ZZFeatureMap+QCNN 1423.51 - 378 116
HEE+QCNN 1468.43 - 89 33
TPE+QCNN 1408.37 - 72 16
[…]